# Maximising Histopathology Segmentation using Minimal Labels via Self-Supervision

### v1

@@ 5 Results
Fully supervised models were trained to establish baselines for different label splits, including 100% labels. It was found that the fine-tuned models consistently outperform fixed-feature models, and therefore only fine-tuned results are shown here (fixed-feature results are in A).
Table 2: A comparison of various self-supervised pre-training methods and respective baselines (randomly initialised without any pre-training) for the downstream tasks of UNet, MDS1, and UDAGAN using various splits of labelled data. For UNet, the labels have been used for all stains, while for MDS1 and UDAGAN, the labels for only source (PAS) stain are used. The evaluation is conducted on an independent, unseen test dataset using F1 score. Each F1 score is the average of five different training repetitions (standard deviations are in parentheses). The highest F1 score for each stain, across different label splits, is in italics, while the overall highest F1 score averaged across all stains is in bold.
@@
SimCLR 0.615 (0.015) 0.403 (0.031) 0.594 (0.026) 0.614 (0.028) 0.556 (0.025)
BYOL 0.516 (0.041) 0.363 (0.027) 0.525 (0.047) 0.494 (0.031) 0.474 (0.037)
HR-CS-CO 0.326 (0.025) 0.224 (0.045) 0.359 (0.050) 0.384 (0.035) 0.323 (0.039)
[…]
SimCLR 0.798 (0.005) 0.534 (0.015) 0.767 (0.008) 0.729 (0.016) 0.707 (0.011)
BYOL 0.713 (0.051) 0.538 (0.047) 0.733 (0.032) 0.605 (0.061) 0.647 (0.048)
HR-CS-CO 0.760 (0.028) 0.335 (0.084) 0.773 (0.015) 0.607 (0.044) 0.619 (0.043)
[…]
SimCLR 0.784 (0.026) 0.541 (0.029) 0.752 (0.040) 0.722 (0.016) 0.700 (0.028)
BYOL 0.706 (0.063) 0.541 (0.060) 0.731 (0.084) 0.650 (0.043) 0.657 (0.062)
HR-CS-CO 0.771 (0.037) 0.433 (0.059) 0.804 (0.041) 0.633 (0.033) 0.660 (0.042)
[…]
SimCLR 0.826 (0.033) 0.638 (0.056) 0.836 (0.034) 0.712 (0.030) 0.753 (0.038)
BYOL 0.833 (0.032) 0.632 (0.042) 0.864 (0.028) 0.652 (0.066) 0.745 (0.042)
HR-CS-CO 0.863 (0.017) 0.614 (0.067) 0.878 (0.018) 0.730 (0.040) 0.771 (0.036)
[…]
@@
On average, in the limited label cases, which are equivalent to 5–6 (1%) and 26–33 (5%) labelled glomeruli per stain, the fine-tuned UNet models significantly outperform the respective baseline UNet models (see last column). This outperformance is not uniform over all stains however, notably Sirius Red and CD34 with 5% labels do benefit from pre-training but not as considerably as the other stains. For some stains, it can be observed that pre-training with 100% labels can even outperform the baseline fully supervised models, however, the benefits are not evident when averaging over all stains. As our goal is to find a labelling level that minimises labelling effort while maximising performance, 5% labels offers a good balance between the two (10% giving only a small increase in performance, while 1% a considerable drop). At this level of labelling, a 11% drop in performance is observed with BYOL pre-trained UNet in comparison to the fully (100%) supervised model. This highlights that the number of labels required for training can be reduced by 95%. If SSL had not been used in this case, a $26.9\%$ drop in performance would have been observed ($5^{\text{th}}$ row, last column of Table 5).
In MDS1 multi-stain segmentation, the same pattern can be observed. Using 1% and 5% labels (but in this case only from the source, PAS, stain) results in a considerable average performance increase over the baseline models. Focusing on 5% labels, SimCLR pre-training enables MDS1 to achieve an average F1 score of $0.707$, which is only 8.2% lower than the 100% supervised MDS1 baseline ($0.789$), while reducing the labelling requirement by 95%. Moreover, this is only 5% lower than the best average UNet single-stain performance with pre-training, which requires labels for all stains, whereas MDS1 requires them for only the source stain.
This trend continues in the stain invariant UDAGAN model’s results, where on average pre-training and fine-tuning with 1% and 5% labels (again, for only the source stain) considerably outperforms the baselines in all stains. HR-CS-CO pre-training is not evaluated as UDAGAN is a stain-invariant single-model multi-stain segmentation approach and HR-CS-CO is trained separately for each stain. In this case, we observe a 10.6% performance drop when fine-tuning with 5% labels (and pre-training with BYOL) compared to the 100% supervised baselines. If the model had been trained in a fully supervised manner with this amount of labels, a $26.1\%$ drop would have been observed, thus fine-tuning is able to minimise the impact of the lack of labels.
@@
MDS1
SimCLR
Refer to caption Refer to caption Refer to caption Refer to caption
BYOL
Refer to caption Refer to caption Refer to caption Refer to caption
HR-CS-CO
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Fig. 9: Visual comparison between predicted glomeruli segmentation maps and real ground-truths for each test stain using fine-tuned UNet, MDS1, and UDAGAN models (trained with 5% labels).
Table 3: Downstream task performance with 5% training labels, without a validation set. UNet, 5% labels are used for all stains, MDS1 and UDAGAN, 5% labels are used for only source, PAS, stain. The evaluation is conducted on test set. Each F1 score is the average of five different training repetitions (standard deviations in parentheses). Highest F1 score for each stain is in italics, overall highest F1 score averaged across all stains is in bold.
@@
Refer to caption Refer to caption
(a) (b)
5.1 Omitting Validation Data
As shown above, a balance between minimising labels and maximising performance is achieved using 5% labels. Nevertheless, when training the final models, the results were obtained using a fully labelled validation set. Therefore, Table 5 evaluates whether the validation set is necessary or whether this labelling requirement can also be reduced. It is shown that in many cases, the performance without a validation set outperforms that obtained when using a labelled validation set. This is explained by the fact that in the dataset used, there is a lower domain shift (measured by following [30]) between the train and test set distributions, which is $0.0655$ (averaged across all stains), compared to the train and validation set distributions, $0.1857$. This allows the models trained without validation data to outperform (on the test data) those selected using the validation loss. Although, this behaviour is specific to datasets with the above-mentioned characteristic, it only affects the difference in performance between the two experimental settings and not the findings themselves. Let us imagine that there were a lower domain shift between the validation and training sets, in this case removing the validation set would only eliminate the increase in performance observed here. It therefore does not invalidate the findings presented herein, that the validation set can be removed to further minimise labelling requirements.
@@ 6 Discussion
The previous section showed the effectiveness of SSL in combating a lack of labelled segmentation data in histopathology, approaching fully-supervised performance (e.g. with BYOL pre-training) in both single-stain UNet and multi-stain UDAGAN models (e.g. with $\sim$30 labels per training stain).
We can observe however that not all self-supervised learning approaches are equal. When fewer labels are available (1% and 5%), general computer vision (CV) approaches such as SimCLR and BYOL perform best. Even though HR-CS-CO is specifically designed for histopathology, it only becomes competitive and/or outperforms the CV approaches when provided with moderate (10%) to larger amounts of labelled data. It is particularly successful when applied to the CD68 stain, outperforming even the baseline models. CD68 is an immunohistochemical stain in which haematoxlyn highlights the main structural component and specific immune cells are highlighted in brown. It is therefore particularly suited to an approach such as HR-CS-CO. There are many other similar immunohistochemical stainings, and more complicated double stainings (e.g. CD3-CD68, CD3-CD163, CD3-CD206, etc [31]) that should be suitable for such an approach (including the H&E stain CS-CO was originally developed for). In some of the other stains used in this study (e.g. Sirius Red), it appears that the superposition of staining components (and weak haematoxyln staining) prevents the haematoxlyn channel from being efficiently extracted, limiting the effectiveness of HR-CS-CO (the difficulty of extracting this component from Sirius Red has been previously noted in the literature [1]).
[…]
@@
Finally, the benefits of self-supervised pre-training are not just restricted to limited label situations. This study has shown that the performance of fully-supervised stain-invariant models such as UDAGAN can be improved—pre-training the UDAGAN model before fully-supervised training lead to a 2.6% increase in F1 score. This offers a new SOTA performance in stain-invariant glomeruli segmentation without any architectural nor labelling changes.
Moving away from renal histopathology, these results are consistent with other histopathology studies found in the literature and extend upon existing efforts to reduce the need for extensive manual annotations. Particularly, Prakash et al. [8] showed that for nuclei segmentation in the Broad Bioimage Benchmark Collection dataset, a self-supervised fine-tuned UNet using only 5% labels (32 images) demonstrated only 3% reduction in IoU score compared to a full supervised UNet. Similarly, Punn et al. [33] reported that a self-supervised fine-tuned UNet using 20% labels (134 images) for nuclei segmentation on the Kaggle Datascience Bowl Challenge 2018 dataset only lost 5.1% in F1 score compared to a fully supervised UNet. Combined with the results presented herein, these demonstrate minimal performance degradation despite significant reductions in label requirements. The findings presented herein, however, go further. Not only do they show the benefit of integrating pre-training into fully-supervised approaches, but also into multi-stain segmentation strategies and removing the need for labelled validation datasets. This more than reduces the labelling requirement to the source stain (a reduction of at least $n$ times, where $n$ is the number of stains to be segmented).
@@ 7 Conclusions
This article has shown how to significantly reduce the need for labelled data ($>$95%) in histopathology image segmentation. To achieve this, self-supervised pre-training techniques— SimCLR, BYOL, and a novel histopathological SSL approach, HR-CS-CO—were used to learn general features from unlabelled data. These features were then fine-tuned for single stain and multi-stain segmentation tasks using UNet, MDS1, and UDAGAN models, making them robust to training scenarios with limited labels.
These approaches demonstrated consistently superior performance compared to their respective baselines, and were able to approach the performance of fully supervised models. These findings underscore the potential and significance of incorporating these advanced learning techniques in histopathology. The results also demonstrated that self-supervised learning combined with fine-tuning is most effective without a validation set, further reducing the labelling requirement. However, some methods, such as SimCLR, are more susceptible to domain shifts and may benefit from some labelled validation data to ensure generalisation.
Furthermore, this study advanced the recent trend in histopathology towards creating multi-stain segmentation models by demonstrating that it is possible to train a stain-invariant segmentation model with as few as $\sim$30 labelled positive patches from one stain. This model closely matches the performance of a fully supervised UNet trained with $\sim$3000 positive patches.
@@ Acknowledgements
Appendix A Results with Fixed Representations

### v2

@@ 5 Results
Fully supervised models were trained to establish baselines for different label splits, including 100% labels. It was found that the fine-tuned models consistently outperform fixed-feature models, and therefore only fine-tuned results are shown here (fixed-feature results are in B.1).
Table 2: A comparison of various self-supervised pre-training methods and respective baselines (randomly initialised without any pre-training) for the downstream tasks of UNet and UDAGAN using various splits of labelled data. For UNet, the labels have been used for all stains, while for UDAGAN, the labels for only source (PAS) stain are used. The evaluation is conducted on an independent, unseen test dataset using F1 score. Each F1 score is the average of five different training repetitions (standard deviations are in parentheses). The highest F1 score for each stain, across different label splits, is in italics, while the overall highest F1 score averaged across all stains is in bold.
@@
On average, in the limited label cases, which are equivalent to 5–6 (1%) and 26–33 (5%) labelled glomeruli per stain, the fine-tuned UNet models significantly outperform the respective baseline UNet models (see last column). This outperformance is not uniform over all stains, however, notably Sirius Red and CD34 with 5% labels do benefit from pre-training but not as considerably as the other stains. For some stains, it can be observed that pre-training with 100% labels can even outperform the baseline fully supervised models, however, the benefits are not evident when averaging over all stains. As our goal is to find a labelling level that minimises labelling effort while maximising performance, 5% labels offers a good balance between the two (10% giving only a small increase in performance, while 1% a considerable drop). At this level of labelling, a 11% drop in performance is observed with BYOL pre-trained UNet in comparison to the fully (100%) supervised model. This highlights that the number of labels required for training can be reduced by 95%. If SSL had not been used in this case, a 26.9% drop in performance would have been observed ($5^{\text{th}}$ row, last column of Table 5).
This trend continues in the stain invariant UDAGAN model’s results, where on average pre-training and fine-tuning with 1% and 5% labels (but in this case only from the source, PAS, stain) considerably outperforms the baselines in all stains. HR-CS-CO pre-training is not evaluated as UDAGAN is a stain-invariant single-model multi-stain segmentation approach and HR-CS-CO is trained separately for each stain. In this case, we observe a 10.6% performance drop when fine-tuning with 5% labels (and pre-training with BYOL) compared to the 100% supervised baselines. If the model had been trained in a fully supervised manner with this amount of labels, a $26.1\%$ drop would have been observed, thus fine-tuning is able to minimise the impact of the lack of labels.
@@
Fig. 7: Visual comparison between predicted glomeruli segmentation maps and real ground-truths for each test stain using fine-tuned UNet and UDAGAN models (trained with 5% labels).
5.1 Omitting Validation Data
As shown above, a balance between minimising labels and maximising performance is achieved using 5% labels. Nevertheless, when training the final models, the results were obtained using a fully labelled validation set. Therefore, Table 5.1 evaluates whether the validation set is necessary or whether this labelling requirement can also be reduced. It is shown that in many cases, the performance without a validation set outperforms that obtained when using a labelled validation set. This is explained by the fact that in the dataset used, there is a lower domain shift (measured by following [31]) between the train and test set distributions, which is $0.0655$ (averaged across all stains), compared to the train and validation set distributions, $0.1857$. This allows the models trained without validation data to outperform (on the test data) those selected using the validation loss. Although, this behaviour is specific to datasets with the above-mentioned characteristic, it only affects the difference in performance between the two experimental settings and not the findings themselves. Let us imagine that there were a lower domain shift between the validation and training sets, in this case removing the validation set would only eliminate the increase in performance observed here. It therefore does not invalidate the findings presented herein, that the validation set can be removed to further minimise labelling requirements.
Refer to caption Refer to caption
(a) (b)
Table 3: Downstream task performance with 5% training labels, without a validation set. UNet, 5% labels are used for all stains; UDAGAN, 5% labels are used for only source, PAS, stain. The evaluation is conducted on test set. Each F1 score is the average of five different training repetitions (standard deviations in parentheses). Highest F1 score for each stain is in italics, overall highest F1 score averaged across all stains is in bold.
@@ 5.1 Omitting Validation Data
6 Generalisation and Reproducibility
To evaluate how well the pre-trained models—SimCLR, BYOL, and the proposed HR-CS-CO—generalise beyond our private dataset, we evaluate their performance on two widely recognised public datasets for kidney glomeruli segmentation. These include: (1) the “HuBMAP — Hacking the Kidney” Kaggle Competition dataset [32]; and the “Kidney Pathology Image Segmentation (KPIs)” Challenge dataset [33]. A short overview of these datasets is provided below:
HuBMAP:
[…]
KPIs:
[…]
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Fig. 9: Illustration of PAS-stained kidney tissue images from three datasets, with each image representing a glomerulus. Representative samples are shown from (a) our in-house dataset, (b) the HuBMAP dataset, and (c) the KPIs dataset.
While both datasets use the same PAS stain as included in our original pre-training set, there exists significant differences in the visual appearance of these stains (see Figure 9). These discrepancies stem from variations in tissue preparation methods, laboratory-specific techniques and the scanning devices [35]. As a result, evaluating our pre-trained models on these datasets offers a rigorous evaluation of their generalisation capabilities beyond the original data. Since both HuBMAP and KPIs contain only single (PAS) stain samples, we used UNet as the downstream segmentation task, following the similar training protocols as detailed in our main UNet experiments (see Section 4.2.2).
[…]
Table 4: Comparison of generalisation performance between self-supervised pre-training methods (trained using 5% labelled data) and their respective baselines (without pre-training) on publicly available kidney glomeruli segmentation datasets. For comparison, the results from fully supervised models (trained with 100% labels) are included, as well as top-performing methods from literature and competition winning models. Performance is evaluated using mean F1 score with standard deviations in parentheses. Best results for each Dataset/Label Split are shown in bold.
{NiceTabular}
ccccc Dataset Label Method Publication Year F1 Score
[…]
SimCLR — 0.736 (0.054)
BYOL — 0.816 (0.080)
HR-CS-CO — 0.814 (0.052)
[…]
9-1100% None (Baseline) — 0.947 (0.002)
SimCLR — 0.922 (0.025)
BYOL — 0.913 (0.022)
HR-CS-CO — 0.920 (0.015)
U-Net SEResNext-101 [32] 2023 0.951 (0.013)
LinkNet [36] 2023 0.942 (–)
SegNeXt [37] 2024 0.868 (–)
CNN-TransXNet [38] 2024 0.828 (–)
V-SAM-VIT-B [39] 2025 0.955 (–)
[…]
SimCLR — 0.751 (0.008)
BYOL — 0.795 (0.016)
HR-CS-CO — 0.651 (0.005)
[…]
6-1100% None (Baseline) — 0.786 (0.023)
SimCLR — 0.848 (0.016)
BYOL — 0.837 (0.014)
HR-CS-CO — 0.794 (0.013)
HoloHisto-4K [41] 2024 0.845 (–)
Mask2Former-Swin-B [40] 2025 0.886 (0.058)
indicates competition-winning method on the respective dataset (HuBMAP: U-Net SEResNext-101; KPIs: Mask2Former-Swin-B), reflecting the leaderboard’s top performance.
Table 6 presents the F1 based segmentation scores achieved using pre-trained models (trained on our private dataset) and fine-tuned on HuBMAP and KPIs with both limited (5%) and fully (100%) labelled data. These results are closely aligned with performance trends observed in the previous section (on our private dataset), providing strong evidence on the generalisation capability of these pre-trained models when applied to new and unseen datasets.
The results show that the pre-trained models have consistent and often substantial gains compared to their baseline counterparts across both datasets. Notably, on the HuBMAP dataset, both BYOL and HR-CS-CO, achieve F1 score of 0.816 and 0.814, respectively, when pre-trained on our dataset and fine-tuned using 5% labels, representing a significant improvement of nearly 10% over the baseline model (0.719). These performance gains are even more pronounced on the KPIs dataset, where BYOL achieves the highest F1 score of 0.795 with just 5% labels. This not only reflects a 12.3% increase over the respective baseline (0.672), but also exceeds the performance of fully supervised baseline (0.786) trained with 100% labels. These findings highlight the effectiveness of these pre-trained models in extracting transferable visual features that facilitate robust performance even with heavily reduced labels. The only notable exception is HR-CS-CO on the KPIs dataset, where its performance (0.651) with 5% labels is slightly below the baseline.
Beyond surpassing the baseline models, the pre-trained models demonstrate a substantial ability to close the performance gap with fully supervised and state-of-the-art benchmarks. For instance, on the HuBMAP dataset, the top performing pre-trained models—BYOL and HR-CS-CO—trained with only 5% labels, achieve over 85% of the F1 score of the competition winning fully supervised method (U-Net SEResNext-101, F1=0.951). These models also show competitive performance relative to other leading fully supervised methods such as LinkNet and SegNeXt. On KPIs, BYOL again stands out with an F1 score of 0.795. This not only surpasses the fully supervised baseline (0.786) but also significantly reduces the performance gap with other fully supervised alternatives (HoloHisto-4K, F1=0.845 and Mask2Former-Swin-B, F1=0.886). Moreover, when trained with 100% labelled data, all pre-trained models (except HR-CS-CO on KPIs) outperform the baseline model with a margin of 5–6%, and their performance closely approaches that of theg top-ranking model (Mask2Former-Swin-B). These findings highlight the effectiveness of these pre-trained models in learning robust and transferable features that enable competitive performance even under extreme annotation scarcity (95% reduced annotations).
To facilitate reproducibility and further research, all our pre-trained models, as well as fine-tuning scripts and detailed instructions, have been made publicly available [^2] https://github.com/zeeshannisar/maximising-kidney-glomeruli-segmentation-using-limited-labels.
@@ 7 Discussion
The previous sections have demonstrated the effectiveness of SSL in combating a lack of labelled segmentation data in histopathology, with pre-trained models approaching fully-supervised performance (e.g. with BYOL) in both single-stain UNet and multi-stain UDAGAN models (e.g. with $\sim$30 annotated images per training stain).
We can observe however that not all self-supervised learning approaches are equal. When fewer labels are available (1% and 5%), general computer vision (CV) approaches such as SimCLR and BYOL perform best. Even though HR-CS-CO is specifically designed for histopathology, it only becomes competitive and/or outperforms the CV approaches when provided with moderate (10%) to larger amounts of labelled data. It is particularly successful when applied to the CD68 stain, outperforming even the baseline models. CD68 is an immunohistochemical stain in which haematoxlyn highlights the main structural component and specific immune cells are highlighted in brown. It is therefore particularly suited to an approach such as HR-CS-CO. There are many other similar immunohistochemical stainings, and more complicated double stainings (e.g. CD3-CD68, CD3-CD163, CD3-CD206, etc [42]) that should be suitable for such an approach (including the H&E stain CS-CO was originally developed for). In stains such as Sirius Red, it appears that the superposition of staining components (and weak haematoxyln staining) prevents the haematoxlyn channel from being efficiently extracted, limiting the effectiveness of HR-CS-CO (the difficulty of extracting this component from Sirius Red has been previously noted in the literature [1]).
The intended application of the pre-trained model naturally dictates the type of SSL that should be used. In UDAGAN, BYOL consistently outperforms SimCLR, especially with highly limited labels, likely due to its robustness to noise during fine-tuning. Unlike SimCLR, which relies on negative pairs, BYOL uses only positive samples and keeps a moving average for regularisation, making it less sensitive to noisy (i.e. translated) data during fine-tuning [43]. It is known that this noise is particularly evident in immunohistochemical stainings such as CD68 and CD34 [4, 31], which is confirmed in this study where the noise degrades the performance of all pre-training methods equally.
@@
Finally, the benefits of self-supervised pre-training are not just restricted to limited label situations. This study has shown that the performance of fully-supervised stain-invariant models such as UDAGAN can be improved, meaning that pre-training the UDAGAN model before fully-supervised training leads to a 2.6% increase in F1 score. This offers a new SOTA performance in stain-invariant glomeruli segmentation without any architectural nor labelling changes.
Moving away from renal histopathology, these results are consistent with other histopathology studies found in the literature and extend upon existing efforts to reduce the need for extensive manual annotations. Particularly, Prakash et al. [12] showed that for nuclei segmentation in the Broad Bioimage Benchmark Collection dataset, a self-supervised fine-tuned UNet using only 5% labels (32 images) demonstrated only 3% reduction in IoU score compared to a full supervised UNet. Similarly, Punn et al. [44] reported that a self-supervised fine-tuned UNet using 20% labels (134 images) for nuclei segmentation on the Kaggle Datascience Bowl Challenge 2018 dataset only lost 5.1% in F1 score compared to a fully supervised UNet. Combined with the results presented herein, these demonstrate minimal performance degradation despite significant reductions in label requirements. The findings presented herein, however, go further. Not only do they show the benefit of integrating pre-training into fully-supervised approaches, but also into multi-stain segmentation strategies and remove the need for labelled validation datasets. This leads to a significant decrease in labelling requirement to the source stain (a reduction of at least $n$ times, where $n$ is the number of stains to be segmented).
@@ 8 Conclusions
This article has shown how to significantly reduce the need for labelled data (>>95%) in histopathology image segmentation. To achieve this, self-supervised pre-training techniques—SimCLR, BYOL, and a novel histopathological SSL approach, HR-CS-CO—were used to learn general features from unlabelled data. These features were then fine-tuned for single stain and multi-stain segmentation tasks using UNet, and UDAGAN models, making them robust to training scenarios with limited labels.
These approaches demonstrated consistently superior performance compared to their respective baselines, and were able to approach the performance of fully supervised models. Moreover, the pre-trained models were generalisable beyond their training datasets. These findings underscore the potential and significance of incorporating these advanced learning techniques in histopathology. The results also demonstrated that self-supervised learning combined with fine-tuning is most effective without a validation set, further reducing the labelling requirement. However, some methods, such as SimCLR, are more susceptible to domain shifts and may benefit from some labelled validation data to ensure generalisation.
Furthermore, this study advanced the recent trend in histopathology towards creating multi-stain segmentation models by demonstrating that it is possible to train a stain-invariant segmentation model with as few as 5% labels from only one stain. This model closely matches the performance of a fully supervised UNet trained with $N\times 100\%$ labels, where $N$ is the number of stains.
@@ Acknowledgements
Appendix A Methods
A.1 Self-supervised Pre-training
A.1.1 SimCLR
Fig. 10: Overview of the SimCLR architecture inspired by [8, 16].
SimCLR [8] learns representations by maximising the agreement between two augmented views of the same image via contrastive loss in the latent space. As shown in Fig. 10, the framework starts with a probabilistic data augmentation module $f_{aug}$ that generates two positively correlated views, $x_{i}$ and $x_{j}$, of a given data sample $x$. A set of base augmentations [8] are adopted (using the Albumentations library [46] and parameters from [15]), including random cropping and resizing with a large scale range of (0.21.00.2-1.0), flipping, grey-scale, Gaussian blur, and random colour distortions. Based on the findings of Stacke et al. [15], two additional augmentations were used, grid distort and grid shuffle, which have demonstrated their effectiveness for histopathology related applications. The augmented views, $x_{i}$ and $x_{j}$ are then transformed into their corresponding representations, $h_{i}$ and $h_{j}$ by employing a convolutional neural network (CNN) base encoder $f_{\theta}$, where $\theta$ is the weight parameters.
where $\tau$ is the temperature parameter that weights different samples and facilitates learning from hard negative samples and $\mathds{1}$ is the indicator function. The term $sim(z_{i},z_{j})=z_{i}^{\top}z_{j}/\|z_{i}\|\|z_{j}\|$ represents the dot product between $\ell_{2}$ normalised $z_{i}$ and $z_{j}$, which corresponds to the cosine similarity. This loss functions aims to maximise the agreement between positive pairs of augmented images, while minimising it for other images in the same batch (negative pairs). In each training step with a batchsize of $2N$, each augmented image has one positive and 2(N1)2(N-1) negative pairs.
A.1.2 BYOL
Fig. 11: Overview of BYOL architecture inspired by Grill et al. [9].
$\xi\leftarrow\tau\xi+(1-\tau)\theta.$ (4)
The loss $\mathcal{L}_{\theta,\xi}$ is made symmetrical by separately feeding $x_{j}$ to the Online network and $x_{i}$ to the Target network. This allows the computation of another loss function $\tilde{\mathcal{L}}_{\theta,\xi}$. During each training step, a stochastic optimisation step is performed to minimise $\mathcal{L}^{BYOL}_{\theta,\xi}=\mathcal{L}_{\theta,\xi}+\tilde{\mathcal{L}}_{\theta,\xi}$ with respect to $\theta$ only, while $\xi$ remains unaffected by applying a stop-gradient ($sg$) as shown in Fig. 11.
Appendix B Supplementary Results
B.1 Fixed Representations